OpenStudy (anonymous):

What is the slope of the line 4x + 6y = 12? A.2/3 B.2 C.-2/3 D.6

OpenStudy (anonymous):

Just manipulate it so the slope jumps out at you.\[4x + 6y = 12 \implies y=-\frac{2}{3}x+2 \implies m=-\frac{2}{3}\]

OpenStudy (anonymous):

oh ok thank you!!!

OpenStudy (anonymous):

trick. Always first write the equation in standard slope form: y=....